Suppose that you are a monopolist in the market of a specific video game. Your inverse demand curve and cost function are the following:

P = 80 - (Q/2)

TC(Q) = 400 + [(Q^2)/2]

The MC (Marginal Cost) is:

Select one:

a. Q^2

b. 0

c. Q

d. 2Q

Question

2

Question text

Suppose that you are a monopolist in the market of a specific video game. Your inverse demand curve and cost function are the following:

P = 80 - (Q/2)

TC(Q) = 400 + [(Q^2)/2]

The MR (Marginal Revenue) is:

Select one:

a. 2Q

b. Q

c. 80

d. 80 - Q

Question

3

Question text

Suppose that you are a monopolist in the market of a specific video game. Your inverse demand curve and cost function are the following:

P = 80 - (Q/2)

TC(Q) = 400 + [(Q^2)/2]

The Equilibrium Quantity Q is:

Select one:

a. 40

b. 80

c. 200

d. 60

Question

4

Question text

Suppose that you are a monopolist in the market of a specific video game. Your inverse demand curve and cost function are the following:

P = 80 - (Q/2)

TC(Q) = 400 + [(Q^2)/2]

Total Profit is:

Select one:

a. $0

b. $1,200

c. $2,400

d. $1,000

Question 5

Question text

Suppose that you are a monopolist in the market of a specific video game. Your inverse demand curve and cost function are the following:

P = 80 - (Q/2)

TC(Q) = 400 + [(Q^2)/2]

The Equilibrium Price P is:

Select one:

a. $60

b. $200

c. $40

d. $80
Answer #1

1. C. Total Cost = 400+ Q power 2/2

Differentiating this function we will get marginal cost I-e

MC= 0+ 2Q/2

MC= Q

2. D. Given Price = 80-Q/2

Total Revenue= Price *Quantity

=Q(80-Q/2)= 80Q-Q power 2/2

Differentiation the above equation we get Marginal revenue

MR= 80-2Q/2

MR= 80-Q